OpenStudy (anonymous):

Rewrite each of the following as an exponential expression ( no fraction ): 1.) 1/4 2.) 2/x^3 3.) 1/(2x-3)^3 4.) 3x/(2x)^5

OpenStudy (anonymous):

1.) 1/4 = 1/2² = 2^-2 2.) 2/x^3 = 2x^-3 3.) (2x - 3)^-3 ( if required to expand, do it and then change) 3 3 4.) 3x/(2x)^5 = 3x/32x^5 = ----- x ^(1-5) = ---- x^-4 32 32

OpenStudy (anonymous):

thanx